# Large Language Models Think Too Fast To Explore Effectively

### v1

@@ Abstract
Large Language Models (LLMs) have emerged many intellectual capacities. While numerous benchmarks assess their intelligence, limited attention has been given to their ability to explore—an essential capacity for discovering new information and adapting to novel environments in both natural and artificial systems. The extent to which LLMs can effectively explore, particularly in open-ended tasks, remains unclear. This study investigates whether LLMs can surpass humans in exploration during an open-ended task, using Little Alchemy 2 as a paradigm, where agents combine elements to discover new ones. Results show most LLMs underperform compared to humans, except for the o1 model, with those traditional LLMs relying primarily on uncertainty-driven strategies, unlike humans who balance uncertainty and empowerment. Representational analysis of the models with Sparse Autoencoders (SAE) revealed that uncertainty and choices are represented at earlier transformer blocks, while empowerment values are processed later, causing LLMs to think too fast and make premature decisions, hindering effective exploration. These findings shed light on the limitations of LLM exploration and suggest directions for improving their adaptability.
Keywords Large Language Models  $\cdot$ Exploration  $\cdot$ Empowerment  $\cdot$ Uncertainty  $\cdot$ Reasoning
@@ 1 Introduction
To explore these questions, we adopt the experimental paradigm introduced by Brändle et al. [2], using the video game Little Alchemy 2 (see methods 2.1). In this game, participants aim to create as many elements as possible by combining known elements, a task that closely aligns with the concept of empowerment and offers a robust framework for evaluating open-ended exploration. We then apply regression models to analyze the exploration strategies of humans and LLMs, focusing on both uncertainty-driven and empowerment-based behaviors (see methods 2.3). Finally, we deploy Sparse AutoEncoders (SAE) to probe the latent representations of exploration-related values, providing insights into how LLMs process information and generate exploratory behavior. This study not only enhances our understanding of LLMs’ exploratory abilities but also highlights exploration as a key element for building more adaptive and intelligent AI systems.
2 Methods
@@ 2.1 Task Description: Little Alchemy 2
Little Alchemy 2 involves discovering new elements by combining a predefined set of basic elements: water, fire, earth, and air. These elements serve as the initial inventory, and players (humans or LLMs) attempt to discover new combinations based on deterministic rules(see Figure.1). The total is 720 elements and the elements range across categories including nature, space, animals, plants, food, inventions, technology, science, tools, buildings, lore, and myths. Among these elements, only 3,452 combinations (out of 259,560) can successfully create other elements and therefore, it requires a semantic understanding of the elements (i.e., empowerment) to explore effectively. This framework mimics a creative combinatorial space exploration task, challenging participants to explore patterns to expand their inventory.
@@ 2.2 Experimental Setup
We evaluated the performance of four LLMs: gpt-4o-2024-08-06(gpt-4o, OpenAI Team [13]), o1-2024-12-17(o1, OpenAI Team [14]), Meta-Llama-3.1-8B-Instruct(LLaMA3.1-8B, Llama Team [10]), and Meta-Llama-3.1-70B-Instruct(LLaMA3.1-70B, Llama Team [10]). These models were selected to represent a range of model sizes and architectures, closed source as well as open source, allowing us to analyze impacts from model variations on exploration and discovery. Each model was prompted with game rules, current inventory, and trial history to contextual reasoning (Figure. 1). Their outputs were constrained to valid game actions, in the format of element + element (for complete prompts, see Figure. 7A).
To investigate the impact of randomness on exploration, we varied the sampling temperature across four settings: 0.0, 0.3, 0.7, and 1.0 (o1 is not available to set parameters and defaults as 1), and under each temperature, there are five repetitions for running the experiment. Lower temperatures encourage deterministic outputs, favoring exploitation, while higher temperatures introduce stochasticity, promoting exploration. These settings allowed us to examine the trade-offs between exploring uncertain combinations and leveraging known strategies.
@@ 2.3.1 Empowerment
Empowerment: In the context of Little Alchemy 2, empowerment translates into the players’ intrinsic desire to create elements that offer many new successful combinations.Selecting combinations that maximize future potential (e.g., unlocking paths to more elements). For example, the element “human” in combination with other elements leads to 83 new elements, while “alien” leads to only 1 new element. Thus, the ’human’ element is more empowering than the ’alien’ element. Brändle et al. [2] uses a neural network to predict empowerment because it effectively models the complex combinatorial relationships and potential outcomes within Little Alchemy 2. By leveraging the neural network, the method incorporates multiple factors, including the probabilities of successful combinations, the likelihood of specific results, and the intrinsic empowerment values of resulting elements. This approach ensures an accurate estimation of the empowerment value by capturing both the immediate and future potential combinations.
@@ 2.4 Sparse Autoencoder (SAE) Analysis
Recently, SAE has been proposed to understand the latent representation in language models [3, 6]. To explore how Large Language Models (LLMs) represent cognitive variables such as empowerment and uncertainty in the context of the alchemy game, we used Sparse Auto-Encoders (SAEs) to learn the latent representations of elements within the model. For each layer, we extracted embeddings from each trial’s available element in the choice set and train them in the SAE. Then we correlate each neuron in the hidden layer in SAEs with our target cognitive variables (i.e., choices, uncertainty values, and empowerment values) and find out the most correlated neuron, where we suppose the relevant variable is most strongly represented. This analysis will help us understand how the model represents and processes cognitive information through the transformer blocks. Finally, we conduct an intervention to examine whether ablating the most correlated neuron causally reduces the corresponding exploration strategy employed by the LLM in the task.
@@ 2.5 Most LLMs Performed Worse Than Humans, Except o1
From 29,493 human players, 90% completed fewer than 500 trials. Experiments were set up with 500 trials for the LLMs. On average, LLaMA3.1-8B discovered 9 elements, LLaMA3.1-70B discovered 25 elements, gpt-4o discovered 35 elements, and o1 discouvered 177 elements (Figure.1). In comparison, humans discovered 42 elements on average within 500 trials and 51 elements across all trials.
3 Results
o1 significantly outperformed humans ($t=9.71,p<0.001$), while the other LLMs performed worse (gpt-4o: $t=-5.48,p<0.001$; LLaMA3.1-70B: $t=-6.39,p<0.001$; LLaMA3.1-8B: $t=-28.12,p<0.001$). Performance improved with larger model sizes, with LLaMA3.1-70B outperforming LLaMA3.1-8B ($t=-6.02,p<0.0001$) and gpt-4o slightly surpassing LLaMA3.1-70B ($t=3.27,p=0.003$).
Exploration success declines in later trials as the inventory grows, and it becomes much harder as the probability of success decreases (see Appendix A). Therefore, different exploration strategies could yield very different performances in the task in different phases. Effective strategies in the latter phase rely on understanding latent game structures (empowerment) rather than uncertainty-driven exploration.
Sampling temperatures were manipulated to assess their impact on exploration strategies. Higher temperatures moderately improved performance ($\beta=0.124,z=5.060,p<0.001$).
Behavioral patterns (Figure.2) highlight o1’s superior strategy, achieving more successful outcomes with new combinations and avoiding repetition of failed or already-successful pairings. This underscores o1’s strong exploratory capacity and innovative approach. For other LLMs, Our result shows that for even larger models, as temperature increases, the percentage of choosing existing combinations decreases. This reveals a diminishing number of redundant behaviors among LLMs. More importantly, a majority of these new combinations do not generate new elements, suggesting that the high temperatures only alter the uncertainty-driven exploration strategies but not empowerment, since in larger spaces, only random combinations are not sufficient to perform the task effectively(see Appendix B). This also explains why higher temperatures can moderately improve the model performance but are still distant from human behaviors.
Figure 2: A: Human and LLMs different Temperatures’ Performance.LLM and Human Performance Across Temperatures. For LLMs, we set four temperatures(0, 0.3, 0.7, 1). LLMs (gpt-4o, LLaMA3.1-8B, LLaMA3.1-70B) achieve their best performance at $\text{temperature}=1$. B: Human and LLMs Best Temperatures’ Behaviors. According to whether the combination selected by each trial is repeated, successful, and initial, the behavior of each LLM trial is divided into 5 categories. Compare the temperature at which LLM performs best with humans and o1 behavior. C: LLM Inventory Performance Relative to Human Percentiles.
@@ 3.1 LLMs Primarily Use Uncertainty-driven Strategies but Not Empowerment
To examine the exact strategies that the models are using, we calculated uncertainty and empowerment values for each element (see methods 2.3). Then, based on each LLM’s choices of combinations, each of the elements is encoded as ‘chosen’ or ‘not chosen’. We also used random sampling to balance the proportion of chosen elements and not-chosen elements. Then we use trial numbers, uncertainty values, and empowerment values to predict whether an element is chosen or not. Notably, we used a linear mixed effect model with the consideration of random slope on each of these variables, so that we could get individual estimates of each sample.
Most LLMs show near-zero empowerment weights, significantly lower than humans (Figure.3, left panel). This suggests LLMs rarely use empowerment for decision-making. In contrast, o1 demonstrates the highest empowerment weights, surpassing humans, indicating a human-like strategy of focusing on actions that expand future possibilities.
@@
Figure 4: A: SAE Correlation Analysis. Maximum correlation of uncertainty values across layers, peaking at layer 2. Maximum correlation of empowerment values across layers, peaking at layer 72. Maximum correlation of choices across layers, peaking at layer 1. B: LLaMA3.1-70B Intervention Regression Results. The regression estimates for empowerment, and uncertainty under the original condition(LLaMA3.1-70B, temperature = 1), empowerment intervention (set to 0), and uncertainty intervention (set to 0). C: LLaMA3.1-70B Average Inventory of Interventions. Uncertainty intervention significantly reduces the average inventory, indicating its essential role in model performance.
@@ 3.2 Uncertainty and Choices are Processed Much Earlier Than Empowerment in LLMs
This unbalanced strategy used in traditional LLMs makes us wonder why LLMs could not use empowerment in the game. Theoretically, LLMs should be able to represent the semantic meaning of these elements. Are they really representing such information as empowerment but not using it or do they lack of ability to understand empowerment? To investigate this question, we employed Sparse Auto-Encoders(SAE)(see methods 2.4 to decompose the latent representation of elements in LLMs to figure out whether both empowerment and uncertainty are properly represented during the computation.
Our results suggest that, in LLaMA3.1-70B, the uncertainty value is highly correlated at layer 2($r=0.73$, Figure.4A). This suggests that in LLaMA3.1-70B, the uncertainty value is strongly represented in the hidden states. We also discover a moderate correlation with empowerment value at layer 72 ($r=0.55$, Figure.4A), indicating LLaMA3.1-70B also represents empowerment values in the middle layer. Despite both values being represented in the hidden states of the model, we found that when we run logistic regression models for each neuron to predict the choices, the highest beta weights also occur at layer 1($beta=1.08$, Figure.4A), aligning with the representation of uncertainty values. This explains why the model mainly deploys uncertainty-driven exploration strategies but not empowerment. Interestingly, both choices and uncertainty values are strongly represented at the begining layers, which may indicate that the model already decides before processing ‘empowerment’ values of the elements in later layers.
We additionally intervene in those most correlated neurons to investigate casual relationships between them and model behaviors. We ablated the most correlated neuron (zeroing the latent activation) identified by uncertainty values and empowerment values in the experiment with other identical settings. The regression analysis of intervened model behavior suggests after ablating the empowerment neuron in the SAE, the model’s empowerment strategy use is even smaller (Figure.4B). This establishes a casual relationship that the neuron we identified through SAE can control the model’s empowerment strategy. On the other hand, when we ablated the uncertainty value neuron, the model’s performance catastrophically dropped, with most of the trials invalid (Figure.4C). The valid data are not even sufficient for regression analysis. This suggests that the earlier layer of uncertainty is very sensitive in the in-context learning for understanding the contexts of the task and history, showing a fundamental role in this task. The ablation results further confirm the role of these two representations in the model. Notably, beyond ablation (simply zeroing out the activation) of the neurons, we also tried other intervention factors to see whether they would help improve model performance through a ’neuroscience’ approach. Our results showed that both slightly weakening and strengthening neuron activity will hurt the model performance (see appendix C.3). This could suggest a deeper issue that the current infrastructure of LLMs may fail to do such open-ended exploration tasks.

### v2

@@ Abstract
Large Language Models (LLMs) have emerged with many intellectual capacities. While numerous benchmarks assess their intelligence, limited attention has been given to their ability to explore—an essential capacity for discovering new information and adapting to novel environments in both natural and artificial systems. The extent to which LLMs can effectively explore, particularly in open-ended tasks, remains unclear. This study investigates whether LLMs can surpass humans in exploration during an open-ended task, using Little Alchemy 2 as a paradigm, where agents combine elements to discover new ones. Results show most LLMs underperform compared to humans, except for the o1 model, with those traditional LLMs relying primarily on uncertainty-driven strategies, unlike humans who balance uncertainty and empowerment. Results indicate that traditional reasoning-focused LLMs, such as GPT-4o, exhibit a significantly faster and less detailed reasoning process, limiting their exploratory performance. In contrast, the DeepSeek reasoning model demonstrates prolonged, iterative thought processes marked by repetitive analysis of combinations and past trials, reflecting a more thorough and human-like exploration strategy. Representational analysis of the models with Sparse Autoencoders (SAE) revealed that uncertainty and choices are represented at earlier transformer blocks, while empowerment values are processed later, causing LLMs to think too fast and make premature decisions, hindering effective exploration. These findings shed light on the limitations of LLM exploration and suggest directions for improving their adaptability.
@@ 1 Introduction
To explore these questions, we adopt the experimental paradigm of Brändle et al. [5], using the video game Little Alchemy 2 (see methods 3.1). In this game, participants aim to create as many elements as possible by combining known elements, a task that closely aligns with the concept of empowerment and offers a robust framework for evaluating open-ended exploration. We then apply regression models to analyze the exploration strategies of humans and LLMs, focusing on both uncertainty-driven and empowerment-based behaviors (see methods 3.3). Finally, we deploy Sparse AutoEncoders (SAE) to probe the latent representations of exploration-related values, providing insights into how LLMs process information and generate exploratory behavior. This study not only enhances our understanding of LLMs’ exploratory abilities but also highlights exploration as a key element for building more adaptive and intelligent AI systems.
2 Related Work
Exploration in Reinforcement Learning and Human Learning
Exploration, as a crucial intellectual capacity in both natural and artificial agents, has been extensively investigated in both reinforcement learning and human learning. In reinforcement learning, decades of research have led to the development of principled exploration mechanisms, including optimism in the face of uncertainty [4], posterior sampling [24], and intrinsic motivation based on novelty or prediction error [25, 7]. More recent advances emphasize structured and scalable exploration using ensemble-based uncertainty estimates [23], world-model disagreement [26], and offline unsupervised skill discovery [11]. These algorithmic innovations not only aim to solve high-dimensional, sparse-reward problems but also offer testable analogues to human strategies such as random exploration, goal-directed novelty seeking, and strategic uncertainty reduction.
In human learning, cognitive scientists primarily focus on the neurocomputational mechanisms underlying exploratory behavior. Through computational modeling, researchers have identified exploration strategies commonly used by humans, such as random and directed exploration [37, 15, 35], as well as empowerment [5]. Numerous studies have shown that humans excel at balancing exploitation and exploration [28, 12], and at employing diverse exploratory strategies [36]. Using neuroimaging and non-invasive brain stimulation techniques, neuroscientists have revealed the neural correlates and causal mechanisms of exploration behaviors, strategies, and task representations in regions such as the frontopolar cortex and intraparietal sulcus [8, 41]. These algorithmic and implementational discoveries help explain how and why humans explore effectively.
Understanding LLM Cognition
Recent research has examined the cognitive capacities of large language models (LLMs), covering domains such as decision-making [3], theory of mind [29], analogical reasoning [33], and temporal-difference learning [10]. However, most of these studies focus on evaluating task performance across model variants, offering limited insight into the internal computational mechanisms behind success or failure. Notably, Demircan et al. [10] applied sparse autoencoders (SAEs) [19, 6] to investigate value-state representations in TD-learning, demonstrating both correlational and causal links between activations and learned value structures—an approach that parallels techniques in human neuroscience. Yet, despite growing interest in LLM cognition, the capacity of LLMs to explore remains underexplored. Existing studies on LLM exploration are largely confined to bandit tasks [20, 17], leaving open the question of whether—and why—LLMs fail to explore effectively in open-ended environments.
Agents in Text-Based Games
In a parallel line of research, several works have explored the use of language models as agents in text-based games [1, 2, 40]. In addition, recent studies have deployed LLMs in open-ended environments such as Minecraft [32, 13, 14]. However, these efforts primarily aim to optimize task performance and often lack deeper investigation into the algorithmic and implementational mechanisms underlying the exploration capabilities of LLMs in such environments.
3 Method
@@ 3.1 Task Description: Little Alchemy 2
Little Alchemy 2 involves discovering new elements by combining a predefined set of basic elements: water, fire, earth, and air. These elements serve as the initial inventory, and players (humans or LLMs) attempt to discover new combinations based on deterministic rules (see Figure 1). The total is 720 elements and the elements range across categories including nature, space, animals, plants, food, inventions, technology, science, tools, buildings, lore, and myths. Among these elements, only 3,452 combinations (out of 259,560) can successfully create other elements and therefore, it requires a semantic understanding of the elements (i.e., empowerment) to explore effectively. This framework mimics a creative combinatorial space exploration task, challenging participants to explore patterns to expand their inventory.
@@ 3.2 Experimental Setup
We evaluated the performance of five LLMs: gpt-4o-2024-08-06 (GPT-4o [21]), o1-2024-12-17 (o1 [22]), Meta-Llama-3.1-8B-Instruct (LLaMA3.1-8B [18]), Meta-Llama-3.1-70B-Instruct (LLaMA3.1-70B [18]), and DeepSeek-R1 [9]. These models were selected to represent a range of model sizes and architectures, closed source as well as open source, allowing us to analyze impacts from model variations on exploration and discovery. Each model was prompted with game rules, current inventory, and trial history to contextual reasoning (Figure 1). Their outputs were constrained to valid game actions, in the format of element + element (for complete prompts, see Figure 8A).
To investigate the impact of randomness on exploration, we varied the sampling temperature across four settings: 0.0, 0.3, 0.7, and 1.0 (o1 is not available to set parameters and defaults as 1), and under each temperature, there are five repetitions for running the experiment. Lower temperatures encourage deterministic outputs, favoring exploitation, while higher temperatures introduce stochasticity, promoting exploration. These settings allowed us to examine the trade-offs between exploring uncertain combinations and leveraging known strategies. OpenAI and Deepseek-R1 models experiment are conducted via API calls. LLaMA3.1-70B and LLaMA3.1-8B are launched in computing cluster, with an NVIDIA A100-80GB running 13 hours and 2 hours to complete all the experiments (with all the settings and repetitions).
@@ 3.3.1 Empowerment
Empowerment: In the context of Little Alchemy 2, empowerment translates into the players’ intrinsic desire to create elements that offer many new successful combinations. Selecting combinations that maximize future potential (e.g., unlocking paths to more elements). For example, the element “human” in combination with other elements leads to 83 new elements, while “alien” leads to only 1 new element. Thus, the “human” element is more empowering than the “alien” element. Brändle et al. [5] uses a neural network to predict empowerment because it effectively models the complex combinatorial relationships and potential outcomes within Little Alchemy 2. By leveraging the neural network, the method incorporates multiple factors, including the probabilities of successful combinations, the likelihood of specific results, and the intrinsic empowerment values of resulting elements. This approach ensures an accurate estimation of the empowerment value by capturing both the immediate and future potential combinations.
@@ 3.4 Statistical Analysis
3.5 Thought Analysis
To investigate the differences in reasoning processes between traditional LLMs and specialized reasoning models, we compared GPT-4o and DeepSeek-R1. For GPT-4o, reasoning traces were elicited using Chain-of-Thought (CoT) prompting and extracted from its responses. In contrast, DeepSeek-R1 natively generates reasoning tokens without explicit prompting.
We analyzed the reasoning traces by first segmenting them into individual sentences. Each sentence was then annotated with one of seven predefined reasoning labels: state goals, check current inventory, past trial analysis, element property reasoning, combination analysis, outcome prediction, and final choice. We employed GPT-4.1 as an automated classifier to assign labels to each sentence.
To reduce redundancy and better capture the structure of coherent reasoning segments, consecutive sentences with the same label were merged into a single labeled unit. This allowed for a more interpretable and concise representation of the reasoning flow.
@@ 3.6 Sparse Autoencoder (SAE) Analysis
Recently, SAE has been proposed to understand the latent representation in language models [6, 10]. To explore how Large Language Models (LLMs) represent cognitive variables such as empowerment and uncertainty in the context of the alchemy game, we used Sparse Auto-Encoders (SAEs) to learn the latent representations of elements within the model. For each layer, we extracted embeddings from each trial’s available element in the choice set and train them in the SAE. Then we correlate each neuron in the hidden layer in SAEs with our target cognitive variables (i.e., choices, uncertainty values, and empowerment values) and find out the most correlated neuron, where we suppose the relevant variable is most strongly represented. This analysis will help us understand how the model represents and processes cognitive information through the transformer blocks. Finally, we conduct an intervention to examine whether ablating the most correlated neuron causally reduces the corresponding exploration strategy employed by the LLM in the task. Training and analyzing SAEs across all the layers takes 42 hours with an NVIDIA V100 GPU.
4 Results
@@ 4.1 Most LLMs Performed Worse Than Humans, Except o1
From 29,493 human players, 90% completed fewer than 500 trials. Experiments were set up with 500 trials for the LLMs. On average, LLaMA3.1-8B discovered 9 elements, LLaMA3.1-70B discovered 25 elements, GPT-4o discovered 35 elements, DeepSeek-R1 discovered 85 elements, and o1 discouvered 177 elements (Figure 1). In comparison, humans discovered 42 elements on average within 500 trials and 51 elements across all trials.
o1 and DeepSeek-R1 significantly outperformed humans (o1: $t=9.71,p<0.001$; DeepSeek-R1: $t=3.40,p<0.027$), while the other LLMs performed worse (GPT-4o: $t=-5.48,p<0.001$; LLaMA3.1-70B: $t=-6.39,p<0.001$; LLaMA3.1-8B: $t=-28.12,p<0.001$). Performance improved with larger model sizes, with LLaMA3.1-70B outperforming LLaMA3.1-8B ($t=-6.02,p<0.0001$) and GPT-4o slightly surpassing LLaMA3.1-70B ($t=3.27,p=0.003$).
Exploration success declines in later trials as the inventory grows, and it becomes much harder as the probability of success decreases (see Appendix A). Therefore, different exploration strategies could yield very different performances in the task in different phases. Effective strategies in the latter phase rely on understanding latent game structures (empowerment) rather than uncertainty-driven exploration. Sampling temperatures were manipulated to assess their impact on exploration strategies. Higher temperatures moderately improved performance ($\beta=0.124,z=5.060,p<0.001$).
Behavioral patterns (Figure 2) highlight o1’s superior strategy, achieving more successful outcomes with new combinations and avoiding repetition of failed or already-successful pairings (Figure 10A). This underscores o1’s strong exploratory capacity and innovative approach. For other LLMs, our result shows that for even larger models, as temperature increases, the percentage of choosing existing combinations decreases (Figure 7). This reveals a diminishing number of redundant behaviors among LLMs. More importantly, a majority of these new combinations do not generate new elements, suggesting that the high temperatures only alter the uncertainty-driven exploration strategies but not empowerment, since in larger spaces, only random combinations are not sufficient to perform the task effectively (see Appendix B). This also explains why higher temperatures can moderately improve the model performance but are still distant from human behaviors.
Figure 2: Human and LLMs different Temperatures’ Performance. LLM and Human Performance Across Temperatures. For LLMs, we set four temperatures (0, 0.3, 0.7, 1). LLMs (GPT-4o, LLaMA3.1-8B, LLaMA3.1-70B) achieve their best performance at $\text{temperature}=1$.
@@ 4.2 LLMs Primarily Use Uncertainty-driven Strategies but Not Empowerment
To examine the exact strategies that the models are using, we calculated uncertainty and empowerment values for each element (see methods 3.3). Then, based on each LLM’s choices of combinations, each of the elements is encoded as “chosen” or “not chosen”. We also used random sampling to balance the proportion of chosen elements and not-chosen elements. Then we use trial numbers, uncertainty values, and empowerment values to predict whether an element is chosen or not. In particular, we used a linear mixed-effects model with the consideration of random slope on each of these variables, so that we could get individual estimates of each sample.
Most LLMs show near-zero empowerment weights, significantly lower than humans (Figure 3, left panel). This suggests that LLMs rarely use empowerment for decision making. In contrast, o1 demonstrates the highest empowerment weights, surpassing humans, indicating a human-like strategy of focusing on actions that expand future possibilities.
@@
4.3 Reasoning Depth and Token Usage in DeepSeek-R1 and GPT-4o
To further elucidate the differences in reasoning depth between DeepSeek-R1 and GPT-4o, we analyzed their reasoning processes from in the subset of the first 150 trials, focusing on reasoning length, diversity of reasoning labels used, and total token counts associated with each reasoning step. (Figure 4).
Figure 4: Comparison of reasoning depth and token usage between DeepSeek-R1 and GPT-4o. A: Per-Trial Reasoning Depth. DeepSeek-R1 shows substantially longer reasoning sequences and consistently employs all reasoning labels, while GPT-4o exhibits significantly shorter sequences and fewer reasoning types. B: Per-trial Token Usage by Reasoning Labels. DeepSeek-R1 allocates a significantly higher number of tokens across all reasoning labels, especially emphasizing outcome_prediction and combination_analysis. GPT-4o uses substantially fewer tokens, limiting the depth and breadth of analysis.
The results highlight clear distinctions between the models. DeepSeek-R1 consistently exhibits a significantly longer reasoning length per trial, reflecting extensive deliberation before reaching a decision. In contrast, GPT-4o uses minimal reasoning. Additionally, DeepSeek-R1 systematically leverages all seven reasoning labels, and even shows self-reflection in the form of ’backtracing’ (see Appendix C.2) within each trial, ensuring diverse analytical coverage (Figure 4A). The limited engagement of GPT-4o in reasoning results in a narrower scope of label usage, confirming its tendency toward shallow exploration.
Further analysis of token allocation across reasoning categories revealed substantial differences (Figure 4B). DeepSeek-R1 distributes tokens broadly, with particular emphasis on outcome_prediction, combination_analysis, and past_trial_analysis. This suggests a robust and methodical evaluation of possible outcomes and historical context, consistent with its more effective exploratory performance. GPT-4o allocates far fewer tokens, indicative of limited engagement in deeper analyses, especially in critical reasoning steps such as outcome prediction and combination evaluation.
These findings reinforce that DeepSeek-R1’s superior task performance arises from more deliberate, diverse, and deeper reasoning processes, whereas GPT-4o’s weaker performance stems from superficial reasoning strategies and inadequate exploration depth. This highlights the importance of reasoning thoroughness and token resource allocation for effective exploration in open-ended tasks.
@@ 4.4 Uncertainty and Choices are Processed Much Earlier Than Empowerment in LLMs
This unbalanced strategy used in traditional LLMs makes us wonder why LLMs could not use empowerment in the game. Theoretically, LLMs should be able to represent the semantic meaning of these elements. Are they really representing such information as empowerment but not using it, or do they lack of ability to understand empowerment? To investigate this question, we employed Sparse Auto-Encoders (SAE) (see methods 3.6) to decompose the latent representation of elements in LLMs to figure out whether both empowerment and uncertainty are properly represented during the computation.
Our results suggest that, in LLaMA3.1-70B, the uncertainty value is highly correlated at layer 2 ($r=0.73$, Figure 5). This suggests that in LLaMA3.1-70B, the uncertainty value is strongly represented in the hidden states. We also discover a moderate correlation with empowerment value at layer 72 ($r=0.55$, Figure 5), indicating LLaMA3.1-70B also represents empowerment values in the middle layer. Despite both values being represented in the hidden states of the model, we found that when we run logistic regression models for each neuron to predict the choices, the highest beta weights also occur at layer 1 ($beta=1.08$, Figure 5), aligning with the representation of uncertainty values. This explains why the model mainly deploys uncertainty-driven exploration strategies but not empowerment. Interestingly, both choices and uncertainty values are strongly represented at the beginning layers, which may indicate that the model already makes a decision before processing the “empowerment” values of the elements in later layers. We additionally intervene (see Appendix C.3) in those most correlated neurons to investigate causal relationships between them and model behaviors. Ablating the SAE-identified empowerment neuron selectively suppresses empowerment-guided behavior, while ablating the uncertainty neuron induces catastrophic performance failure, thereby establishing causal control by these two latent representations.
Figure 5: SAE Correlation Analysis. Maximum correlation of uncertainty values across layers, peaking at layer 2. Maximum correlation of empowerment values across layers, peaking at layer 72. Maximum beta weight of choices across layers, peaking at layer 1.